# The Griffiths phase and beyond: A large deviations study of the magnetic susceptibility of the two-dimensional bond-diluted Ising model

### v1

@@ I Introduction
The Ising model is one of the most studied systems in statistical physics and has found applications in many branches of science, see, e.g., Refs. [1, 2, 3, 4, 5, 6]. Its original purpose is to describe ferromagnetism in homogeneous materials with strong uniaxial asymmetry. Often, however, materials are not perfectly homogeneous but exhibit randomly distributed impurities. In such cases, a central question is how these random impurities affect the physical properties of the system in comparison to the idealized pure model [7]. Impurities can be incorporated into the Ising model, for instance, by randomly removing a fraction $1-p$ of the bonds that represent the ferromagnetic interactions between the spins, where $0\leq p\leq 1$. Due to the overall weakened ferromagnetic coupling this leads to a shift in the transition temperature from $T_{\mathrm{f}}:=T_{\mathrm{c}}(p=1)$ in the pure system to a lower temperature $T_{\mathrm{c}}(p),~{}p<1$ in the diluted system. As shown by Harris [8], the critical behavior at the transition differs from that found in the pure system if the specific heat exponent $\alpha$ of the pure system is positive. This is the case for the Ising model in dimensions $d\geq 3$, while $\alpha=0$ in $d=2$. Due to this marginality one expects logarithmic corrections to the leading critical behavior in two dimensions, but according to numerical simulations [9, 10, 11] as well as experiments [12, 13] the critical exponents remain identical to the pure case.
@@
The thermal region between $T_{\mathrm{f}}$ and $T_{\mathrm{c}}(p<1)$ is known as the Griffiths phase [14]. In this regime the order parameter of the ferromagnetic phase transition, the magnetization, remains zero, But arbitrarily large fluctuations of the order parameter become possible. These fluctuations can exist due to large compact structures of ferromagnetic bonds in regions where there are less missing bonds in comparison to the overall average of the system. Within these structures the system is effectively in a ferromagnetic state such that a change of the spin orientation happens coherently, giving rise to large fluctuations in the magnetization. As a consequence the distribution of the second moment of the magnetization, the magnetic susceptibility, is expected to have an exponential tail that extends to infinity, which is an expression of the essential, but weak, Griffiths singularity [15, 16]: Since the magnetic susceptibility characterizes the response of the system to an external magnetic field, the free energy is a non-analytic function of the field throughout the Griffiths phase [14]. Besides these effects on static averages, the Griffiths singularity also plays an important role for the dynamics of the system, leading to a slow-down of the decay of the spin-spin correlation function [15, 17, 18]. The Griffiths singularity does not only occur in the diluted Ising ferromagnet [19] but it may also be observed in other disordered systems such as spin glasses [20]. The analogous quantum mechanical effect is known as the Griffiths-McCoy singularity, and it has been studied theoretically, numerically as well as in experiments [21, 22, 23, 24, 25, 26, 27].
@@ II The Two-Dimensional Bond-Diluted Ising Ferromagnet
where $\bm{S}\in\{-1,1\}^{N}$ denotes a spin configuration, and $\hat{M}=\sum_{\bm{x}}s_{\bm{x}}$ represents the configurational magnetization which couples to the external magnetic field $h$. The Ising spins $s_{\bm{x}}=\pm 1$ are placed at the sites $\bm{x}$ of a two-dimensional square lattice of linear dimension $L$, resulting in $N=L^{2}$ spins in total. The notation $\langle\bm{x},\bm{y}\rangle$ refers to summation over nearest neighbors only. The time-independent, quenched interaction between two spins is represented by the exchange coupling $J_{\bm{x}\bm{y}}$. To incorporate random dilution into the model, the bonds are drawn from a bimodal distribution, such that the probability to obtain a particular bond sample $\bm{J}=\{J_{\bm{xy}}\}$ is given by
$P_{J}(\bm{J})=\prod_{\langle\bm{x},\bm{y}\rangle}p\,\delta[J_{\bm{xy}}-1]+(1-p)\delta[J_{\bm{xy}}].$ (2)
where we use $\delta[x]$, $x\in\mathbb{R}$, as an indicator function that yields one if $x=0$ and zero otherwise. Here, $p$ corresponds to the probability of drawing a ferromagnetic bond with $J_{\bm{x}\bm{y}}=1$, while $1-p$ is the probability of missing bonds with $J_{\bm{x}\bm{y}}=0$. The model is studied in the canonical ensemble at temperature $T$, such that the spin configurations $\bm{S}$ are Boltzmann distributed according to
@@
where we have set the Boltzmann constant $k_{\mathrm{B}}:=1$ for convenience. $Z_{\bm{J}}$ is the partition function for a given bond sample $\bm{J}$. The Gibbs state of the model in the absence of an external magnetic field, $h=0$, depends on two parameters, the bond occupation probability $p$ and the temperature $T$. The order parameter which can be used to determine the phase of the Gibbs state is the first moment of the magnetization per site  [^1] Note that below the ferromagnetic phase transition the symmetry is broken such that there are two Gibbs states, one with positive and one with negative magnetization. To obtain the correct value for the magnetization, the average has to be restricted to one of the Gibbs states. In numerical simulations, however, there is usually no symmetry breaking and thus one often uses the absolute value of the magnetization as an estimator for the magnetization or, alternatively, the largest FKCK cluster, see Sec. III.
@@
where $\hat{m}=\hat{M}/N$. Here, $\langle\cdot\rangle_{S}$ denotes the thermal average with respect to the Boltzmann distribution, Eq. (3), and $[\cdot]_{J}$ is the average with respect to the disorder distribution, Eq. (2). As is clear from the phase diagram of the system shown in Fig. 1, there exists a high temperature paramagnetic phase and a low temperature ferromagnetic phase. The boundary between these phases extends from the bond-percolation threshold $p_{\mathrm{th}}=0.5$ at which the transition temperature is zero, $T_{\mathrm{c}}(0.5)=0$, to that of the pure ferromagnet with $p=1$ and $T_{\mathrm{f}}=T_{\mathrm{c}}(1)=2/\ln(1+\sqrt{2})=2.2691\dots$. In between these limits, the phase boundary can be obtained by arguments based on duality [36, 34] which yield a good estimate for the true curve, consistent with what is found in numerical simulations [37]. The thermal region between the ferromagnetic phase transition of the pure system at $T_{\mathrm{f}}=T_{\mathrm{c}}(1)$ and the ferromagnetic phase transition in the diluted system $T_{\mathrm{c}}(p<1)$ is known as the Griffiths phase. Inside the Griffiths phase the order parameter remains zero but large fluctuations of the magnetization are more likely than in the paramagnetic phase. These fluctuations are visible in the magnetic susceptibility $\chi_{\bm{J}}$ which can be defined from the variance of the magnetization per lattice site for a given bond sample $\bm{J}$,
@@
for $\chi\to\infty$. Here, $A$ is a temperature-dependent positive constant which vanishes at the ferromagnetic phase transition (implying a particularly broad distribution there) and diverges when $T$ approaches $T_{\mathrm{f}}$ (such that the tail disappears in the non-disordered limit). Since the derivation which leads to Eq. (6) includes variational arguments and a number of approximations, Bray concluded that this form may only constitute a lower bound for the true tail. The magnetic susceptibility describes the linear response of the magnetization to an external magnetic field, since $(\partial/\partial h)\langle\hat{m}\rangle_{S}=\chi_{\bm{J}}/T$. Correspondingly, inside the Griffiths phase the free energy is a non-analytic function of the external magnetic field [14].
@@ III Large Deviations Sampling
Here, $Z_{\Theta}$ is a normalization constant that will be determined later, and $f_{\Theta}(Y_{\bm{J}})$ is a bias function which depends on the set of parameters $\Theta$. The bias function has to be chosen such that $\widetilde{P}_{J}(\bm{J};\Theta)$ defines a distribution which has enough weight in the regions that one would like to sample. Following the proposal of Neuhaus and Hager in Ref. [33] for a related problem, we conjecture that in many cases a generalized exponential of the form
@@
with $\Theta=(\theta_{1},\theta_{2},\theta_{3})\in\mathbb{R}^{3}$ serves this purpose well. To explain the effect of parameter $\theta_{1}$, we first set $\theta_{2}=0$ such that $f_{\Theta}$ becomes a simple exponential. Now, if $\theta_{1}<0$ the weight of the bond samples with small $Y_{\bm{J}}$ in comparison to typical values of the unbiased distribution increases, and if $\theta_{1}>0$ bond samples with large $Y_{\bm{J}}$ have more weight. If, on the other hand, we set $\theta_{1}$ to zero, $f_{\Theta}$ has the shape of a Gaussian of width $2/\theta_{2}$ centered around $\theta_{3}$. This means the bias function will increase the weight close to $\theta_{3}$. With an adequate combination of the parameters in $\Theta$ it is possible to create a biased bond distribution with enough weight in regions that lead to the desired values of $Y_{\bm{J}}$. While this recipe is fairly general, we do not exclude the possibility that for some problems other types of bias functions might work better.
@@
To sample from the biased bond distribution $\widetilde{P}_{J}$, the Metropolis-Hastings algorithm is used, which in this particular case works as follows. Suppose that we start with a bond sample $\bm{J}_{\mu}$ with the corresponding value $Y_{\bm{J}_{\mu}}$ of the observable of interest. Now, one generates a candidate bond sample $\bm{J}_{\mu}^{\prime}$ by randomly selecting one or more bonds of $\bm{J}_{\mu}$ and assigning to them new coupling values according to the unbiased distribution (2). In other words, a randomly selected bond is set to $1$ with probability $p$ and to 00 with probability $1-p$. After computing the observable value for the proposed sample, $Y_{\bm{J}_{\mu}^{\prime}}$, the new bond configuration is accepted with probability
@@
Figure 4: The mean magnetic susceptibility $\overline{\chi}$ as a function of temperature $T$ for a fraction $p=0.6$ of ferromagnetic bonds. The values for the different configurations ${\bm{J}}$ are computed according to the estimator of Eq. (13). The grey lines are continuations of the low-temperature cluster estimator into the high-temperature phase. These curves exhibit a peak which shifts towards the critical temperature on increasing the system size, similar to other finite-size definitions of $\chi$. The Griffiths phase extends from the critical temperature of the pure ferromagnet $T_{\mathrm{f}}=2.2691\dots$ to the corresponding temperature of the diluted system $T_{\mathrm{c}}(0.6)=0.9541(10)$.
@@
To estimate the value of the magnetic susceptibility for a given bond configuration it is necessary to compute an average over multiple thermal samples (since the susceptibility is no “configurational” quantity [40]). To this end, we carry out a thermal Monte Carlo simulation utilizing the Swendsen-Wang cluster algorithm [41]. This method is known to perform well for diluted ferromagnets [42], and it is possible to use improved cluster estimators for the susceptibility [43]. Furthermore, the clusters underlying the algorithm provide an interesting geometrical interpretation of the magnetic susceptibility. To elucidate this context, in the following we hence provide a short exposition of the Swendsen-Wang algorithm as well as the corresponding cluster framework. Starting with a given spin configuration, in the Swendsen-Wang algorithm one occupies each bond $J_{\bm{x}\bm{y}}>0$ with probability $p_{\mathrm{FK}}(J_{\bm{x}\bm{y}})=1-\exp(-2/T)$ if the two connected spins are parallel, i.e., if $s_{\bm{x}}s_{\bm{y}}=1$, and $p_{\mathrm{FK}}(J_{\bm{x}\bm{y}})=0$ if $s_{\bm{x}}s_{\bm{y}}=-1$. On the contrary, diluted bonds with $J_{\bm{x}\bm{y}}=0$ are never occupied. Two spin sites which are connected by a path of occupied bonds belong to the same cluster. Clusters which are defined in this way are denoted as FKCK (Fortuin-Kasteleyn–Coniglio-Klein) clusters  [^3] The FKCK clusters are also called FK clusters or CK droplets due to their historical origin [2], see also Ref. [57].. The smallest possible FKCK cluster contains only a single spin site. After constructing the clusters, each of them is randomly assigned an up or down orientation and the spins in each cluster are flipped accordingly. As a result all spins within each cluster have identical sign but the sign of two spins in different clusters may differ. This generates the next spin configuration of the Monte Carlo process. The Swendsen-Wang algorithm is ergodic and satisfies the detailed balance condition with respect to the Boltzmann distribution [29]. The cluster estimator of the magnetic susceptibility is based on the densities $\hat{\rho}_{i}$ of the clusters which are defined as their number of sites divided by $N$. We assume that the indices $i$ are sorted by cluster size such that $\hat{\rho}_{1}$ corresponds to the largest cluster. For $T\geq T_{\mathrm{c}}$ the estimator is given by the average cluster size of all FKCK clusters [43]. For $T<T_{\mathrm{c}}$ it is necessary to subtract the square of the density of the infinite cluster [45, 29, 46, 47],
@@
To numerically estimate $\chi_{\bm{J}}$, we perform a thermal average over $N_{\mathrm{S}}$ cluster configurations. Due to the computational complexity of the problem successive configurations from the Monte Carlo chain are used, such that these individual estimates are correlated. The computational complexity emerges since one has to compute a thermal average each time a new bond sample is proposed. While we need to compute $\chi_{\bm{J}}$ for each bond update, measurements are only recorded for analysis after each sweep consisting of $2N$ bond updates according to the bond Monte Carlo algorithm, see Eq. (10). Figure 2 shows the sampling of $\chi_{\bm{J}}$ for systems of size $L=10$ with $p=0.6$ and $N_{\mathrm{S}}=2000$ samples at temperature $T=2$. The parameters used to generate the data are listed in Table 1. As discussed above, $\theta_{1}<0$ leads to a decrease of $\chi$ and $\theta_{1}>0$ leads to an increase of $\chi_{\bm{J}}$ compared to the unbiased case $\Theta_{u}=(0,0,0)$. For $\Theta_{3}=(0.9,0,0)$ the Monte Carlo chain oscillates between two equilibrium states. The gap between these states increases with system size such that sampling in the intermediate region becomes difficult, and the algorithm is not well suited to study larger systems with $\theta_{2}=\theta_{3}=0$ [31]. This problem can be circumvented by using an appropriate combination of non-zero values for $\theta_{2}$ and $\theta_{3}$; this is illustrated in Fig. 2. The Figure shows the data of 11 parameter sets for a system of size $L=10$. For larger systems one has to increase the number of parameter sets to sample the distribution over a wide range of the support such that computations become more and more time-consuming. To generate the histogram for the largest studied system size $L=128$, see Fig. 13, we used around 320 parameter sets.

### v2

@@ I Introduction
The Ising model is one of the most studied systems in statistical physics and has found applications in many branches of science, see, e.g., Refs. [1, 2, 3, 4, 5, 6, 7]. Its original purpose is to describe ferromagnetism in homogeneous materials with strong uniaxial asymmetry. Often, however, materials are not perfectly homogeneous but exhibit randomly distributed impurities. In such cases, a central question is how these random impurities affect the physical properties of the system in comparison to the idealized pure model [8, 9, 10]. Impurities can be incorporated into the Ising model, for instance, by randomly removing a fraction $1-p$ of the bonds that represent the ferromagnetic interactions between the spins, where $0\leq p\leq 1$. Due to the overall weakened ferromagnetic coupling this leads to a shift in the transition temperature from $T_{\mathrm{f}}:=T_{\mathrm{c}}(p=1)$ in the pure system to a lower temperature $T_{\mathrm{c}}(p),~{}p<1$ in the diluted system. As shown by Harris [11], the critical behavior at the transition differs from that found in the pure system if the specific heat exponent $\alpha$ of the pure system is positive. This is the case for the Ising model in dimensions $d\geq 3$, while $\alpha=0$ in $d=2$. Due to this marginality one expects logarithmic corrections to the leading critical behavior in two dimensions [12, 13, 14, 15, 10], but according to numerical simulations [16, 17, 18, 19, 20] as well as experiments [21, 22] the critical exponents remain identical to the pure case.
@@
The thermal region between $T_{\mathrm{f}}$ and $T_{\mathrm{c}}(p<1)$ is known as the Griffiths phase [23]. In this regime the order parameter of the ferromagnetic phase transition, the magnetization, remains zero, but arbitrarily large fluctuations of the order parameter become possible. These fluctuations can exist due to large compact structures of ferromagnetic bonds in regions where there are less missing bonds in comparison to the overall average of the system. Within these structures the system is effectively in a ferromagnetic state such that a change of the spin orientation happens coherently, giving rise to large fluctuations in the magnetization. As a consequence the distribution of the second moment of the magnetization, the magnetic susceptibility, is expected to have an exponential tail that extends to infinity, which is an expression of the essential, but weak, Griffiths singularity [24, 25]: Since the magnetic susceptibility characterizes the response of the system to an external magnetic field, the free energy is a non-analytic function of the field throughout the Griffiths phase [23]. Besides these effects on static averages, the Griffiths singularity also plays an important role for the dynamics of the system, leading to a slow-down of the decay of the spin-spin correlation function [26, 24, 27, 28]. The Griffiths singularity does not only occur in the diluted Ising ferromagnet [29] but it may also be observed in other disordered systems such as spin glasses [30]. The analogous quantum mechanical effect is known as the Griffiths-McCoy singularity, and it has been studied theoretically, numerically as well as in experiments [31, 32, 33, 34, 35, 36, 37].
@@ II The Two-Dimensional Bond-Diluted Ising Ferromagnet
where $\bm{S}\in\{-1,1\}^{N}$ denotes a spin configuration, and $\hat{M}=\sum_{\bm{x}}s_{\bm{x}}$ represents the magnetization which couples to the external magnetic field $h$. The Ising spins $s_{\bm{x}}=\pm 1$ are placed at the sites $\bm{x}$ of a two-dimensional square lattice of linear dimension $L$, resulting in $N=L^{2}$ spins in total. The notation $\langle\bm{x},\bm{y}\rangle$ refers to summation over nearest neighbors only. The time-independent, quenched interaction between two spins is represented by the exchange coupling $J_{\bm{x}\bm{y}}$. To incorporate random dilution into the model, the bonds are drawn from a bimodal distribution, such that the probability to obtain a particular bond sample $\bm{J}=\{J_{\bm{xy}}\}$ is given by
$P_{J}(\bm{J})=\prod_{\langle\bm{x},\bm{y}\rangle}p\,\delta[J_{\bm{xy}}-1]+(1-p)\delta[J_{\bm{xy}}],$ (2)
where we use $\delta[x]$, $x\in\mathbb{R}$, as an indicator function that yields one if $x=0$ and zero otherwise. Here, $p$ corresponds to the probability of drawing a ferromagnetic bond with $J_{\bm{x}\bm{y}}=1$, while $1-p$ is the probability of missing bonds with $J_{\bm{x}\bm{y}}=0$. The model is studied in the canonical ensemble at temperature $T$, such that the spin configurations $\bm{S}$ are Gibbs-Boltzmann distributed according to
@@
where we have set the Boltzmann constant $k_{\mathrm{B}}:=1$ for convenience. $Z_{\bm{J}}$ is the partition function for a given bond sample $\bm{J}$. The thermodynamic state of the model in the absence of an external magnetic field, $h=0$, depends on two parameters, the bond occupation probability $p$ and the temperature $T$. The order parameter which can be used to determine the phase of the thermodynamic state is the first moment of the magnetization per site  [^1] Note that below the ferromagnetic phase transition the symmetry is broken such that there are two thermodynamic states, one with positive and one with negative magnetization. To obtain the correct value for the magnetization, the average has to be restricted to one of the states. In numerical simulations, however, there is usually no symmetry breaking and thus one often uses the absolute value of the magnetization as an estimator for the magnetization or, alternatively, the largest FKCK cluster, see Sec. III.
@@
where $\hat{m}=\hat{M}/N$. Here, $\langle\cdot\rangle_{S}$ denotes the thermal average with respect to the Gibbs-Boltzmann distribution, Eq. (3), and $[\cdot]_{J}$ is the average with respect to the disorder distribution, Eq. (2). As is clear from the phase diagram of the system shown in Fig. 1, there exists a high temperature paramagnetic phase and a low temperature ferromagnetic phase. The boundary between these phases extends from the bond-percolation threshold $p_{\mathrm{th}}=0.5$ at which the transition temperature is zero, $T_{\mathrm{c}}(0.5)=0$, to that of the pure ferromagnet with $p=1$ and $T_{\mathrm{f}}=T_{\mathrm{c}}(1)=2/\ln(1+\sqrt{2})=2.2691\dots$. In between these limits, the phase boundary can be obtained by arguments based on duality [46, 44] which yield a good estimate for the true curve, consistent with what is found in numerical simulations [47]. The thermal region between the ferromagnetic phase transition of the pure system at $T_{\mathrm{f}}=T_{\mathrm{c}}(1)$ and the ferromagnetic phase transition in the diluted system $T_{\mathrm{c}}(p<1)$ is known as the Griffiths phase. Inside the Griffiths phase the order parameter remains zero but large fluctuations of the magnetization are more likely than in the paramagnetic phase. These fluctuations are visible in the magnetic susceptibility $\chi_{\bm{J}}$ which can be defined from the variance of the magnetization per lattice site for a given bond sample $\bm{J}$,
@@
for $\chi\to\infty$. Here, $A$ is a temperature-dependent positive constant which vanishes at the ferromagnetic phase transition (implying a particularly broad distribution there) and diverges when $T$ approaches $T_{\mathrm{f}}$ (such that the tail disappears). Since the derivation which leads to Eq. (6) includes variational arguments and a number of approximations, Bray concluded that this form may only constitute a lower bound for the true tail. The magnetic susceptibility describes the linear response of the magnetization to an external magnetic field, since $(\partial/\partial h)\langle\hat{m}\rangle_{S}=\chi_{\bm{J}}/T$. Correspondingly, inside the Griffiths phase the free energy is a non-analytic function of the external magnetic field [23].
@@ III Large Deviations Sampling
Here, $Z_{\Theta}$ is a normalization constant that will be determined later, and $f_{\Theta}(Y_{\bm{J}})$ is a bias function which depends on the set of parameters $\Theta$. The bias function has to be chosen such that $\widetilde{P}_{J}(\bm{J};\Theta)$ defines a distribution which has enough weight in the regions that one would like to sample. Following the proposal of Neuhaus and Hager in Ref. [43] to use a sequence of Gaussians centred at successive values of the reaction coordinate in order to bias the probability density in a region of suppressed weight, we propose to use a generalized exponential of the form
@@
with $\Theta=(\theta_{1},\theta_{2},\theta_{3})\in\mathbb{R}^{3}$ to bias the sampling for the purpose of accessing rare events. To understand the effect of this bias in detail, at first set $\theta_{2}=0$ such that $f_{\Theta}$ becomes a simple exponential. Now, if $\theta_{1}<0$ the weight of the bond samples with small $Y_{\bm{J}}$ in comparison to typical values of the unbiased distribution increases, and if $\theta_{1}>0$ bond samples with large $Y_{\bm{J}}$ have more weight. If, on the other hand, we set $\theta_{1}$ to zero, $f_{\Theta}$ has the shape of a Gaussian of width $2/\theta_{2}$ centered around $\theta_{3}$. This means that the bias function will increase the weight close to $\theta_{3}$. With an adequate combination of the parameters in $\Theta$ it is possible to create a biased bond distribution with enough weight in regions that lead to the desired values of $Y_{\bm{J}}$. While this recipe is fairly general, we do not exclude the possibility that for some problems other types of bias functions might work better.
@@
To sample from the biased bond distribution $\widetilde{P}_{J}$, the Metropolis-Hastings algorithm is used, which in this particular case works as follows. Suppose that we start with a bond sample $\bm{J}_{\mu}$ with the corresponding value $Y_{\bm{J}_{\mu}}$ of the observable of interest. Now, one generates a candidate bond sample $\bm{J}_{\mu}^{\prime}$ by randomly selecting one or more bonds of $\bm{J}_{\mu}$ and assigning to them new coupling values according to the unbiased distribution Eq. (2). In other words, a randomly selected bond is set to $1$ with probability $p$ and to 00 with probability $1-p$. After computing the observable value for the proposed sample, $Y_{\bm{J}_{\mu}^{\prime}}$, the new bond configuration is accepted with probability
@@
Figure 4: The mean magnetic susceptibility $\overline{\chi}$ as a function of temperature $T$ for a fraction $p=0.6$ of ferromagnetic bonds. The values for the different configurations ${\bm{J}}$ are computed according to the estimator of Eq. (13). The grey lines are continuations of the low-temperature cluster estimator into the high-temperature phase. These curves exhibit a peak which shifts towards the critical temperature on increasing the system size, similar to other finite-size definitions of $\overline{\chi}$. The Griffiths phase extends from the critical temperature of the pure ferromagnet $T_{\mathrm{f}}=2.2691\dots$ to the corresponding temperature of the diluted system $T_{\mathrm{c}}(0.6)=0.9541(10)$.
@@
To estimate the value of the magnetic susceptibility for a given bond configuration it is necessary to compute an average over multiple thermal samples (since the susceptibility is no “configurational” quantity [50], i.e., it cannot be estimated from a single spin configuraton). To this end, we carry out a thermal Monte Carlo simulation utilizing the Swendsen-Wang cluster algorithm [51]. This method is known to perform well for diluted ferromagnets [52], and it is possible to use improved cluster estimators for the susceptibility [53]. Furthermore, the clusters underlying the algorithm provide an interesting geometrical interpretation of the magnetic susceptibility. To elucidate this context, in the following we hence provide a short exposition of the Swendsen-Wang algorithm as well as the corresponding cluster framework. Starting with a given spin configuration, in the Swendsen-Wang algorithm one occupies each ferromagnetic bond $J_{\bm{x}\bm{y}}=1$ with probability $p_{\mathrm{FK}}(J_{\bm{x}\bm{y}})=1-\exp(-2/T)$ if the two connected spins are parallel, i.e., if $s_{\bm{x}}s_{\bm{y}}=1$, and $p_{\mathrm{FK}}(J_{\bm{x}\bm{y}})=0$ if $s_{\bm{x}}s_{\bm{y}}=-1$. On the contrary, diluted bonds with $J_{\bm{x}\bm{y}}=0$ are never occupied. Two spin sites which are connected by a path of occupied bonds belong to the same cluster. Clusters which are defined in this way are denoted as FKCK (Fortuin-Kasteleyn–Coniglio-Klein) clusters  [^3] The FKCK clusters are also called FK clusters or CK droplets due to their historical origin [2], see also Ref. [69].. The smallest possible FKCK cluster contains only a single spin site. After constructing the clusters, each of them is randomly assigned an up or down orientation and the spins in each cluster are flipped accordingly. As a result all spins within each cluster have identical sign but the sign of two spins in different clusters may differ. This generates the next spin configuration of the Monte Carlo process. The Swendsen-Wang algorithm is ergodic and satisfies the detailed balance condition with respect to the Gibbs-Boltzmann distribution [39]. The cluster estimator of the magnetic susceptibility is based on the densities $\hat{\rho}_{i}$ of the clusters which are defined as their number of sites divided by $N$. We assume that the indices $i$ are sorted by cluster size such that $\hat{\rho}_{1}$ corresponds to the largest cluster. For $T\geq T_{\mathrm{c}}$ the estimator is given by the average cluster size of all FKCK clusters [53]. For $T<T_{\mathrm{c}}$ it is necessary to subtract the square of the density of the infinite cluster [55, 39, 56, 57],
@@
To numerically estimate $\chi_{\bm{J}}$, we perform a thermal average over $N_{\mathrm{S}}$ cluster configurations. Due to the computational complexity of the problem successive configurations from the Monte Carlo chain are used, such that these individual estimates are correlated. The computational complexity emerges since one has to compute a thermal average each time a new bond sample is proposed. While we need to compute $\chi_{\bm{J}}$ for each bond update, measurements are only recorded for analysis after each sweep consisting of $2N$ bond updates according to the bond Monte Carlo algorithm, see Eq. (10). Figure 2 shows the sampling of $\chi_{\bm{J}}$ for systems of size $L=10$ with $p=0.6$ and $N_{\mathrm{S}}=2000$ samples at temperature $T=2$. The parameters used to generate the data are listed in Table 1. As discussed above, $\theta_{1}<0$ leads to a decrease of $\chi$ and $\theta_{1}>0$ leads to an increase of $\chi_{\bm{J}}$ compared to the unbiased case $\Theta_{u}=(0,0,0)$. For $\Theta_{3}=(0.9,0,0)$ the Monte Carlo chain oscillates between two equilibrium states. The gap between these states increases with system size such that sampling in the intermediate region becomes difficult, and the algorithm is not well suited to study larger systems with $\theta_{2}=\theta_{3}=0$ [41]. This problem can be circumvented by using an appropriate combination of non-zero values for $\theta_{2}$ and $\theta_{3}$; this is illustrated in Fig. 2. The Figure shows the data of 11 parameter sets for a system of size $L=10$. For larger systems one has to increase the number of parameter sets to sample the distribution over a wide range of the support such that computations become more and more time-consuming. To generate the histogram for the largest studied system size $L=128$, see Fig. 13, we used around 300 parameter sets. The algorithm for choosing these parameters is to a certain degree heuristic in that the Gaussians are chosen in number and position in such a way as to fill the relevant range of $\chi$ while maintaining sufficient overlap in the histograms of individual simulations that is required for the stitching together of simulations according to Eq. (12).